OpenStudy (anonymous):

Paul plans to put concrete on a rectangular portion of his driveway. The portion is 12 feet long and 6 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Paul needs is $108.89. Which of the following is closest to the width of the portion of the driveway on which Paul plans to put concrete? [1 foot = 12inches; 1 yard = 3 feet] 3 feet 5 feet 8 feet 10 feet

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

@donka74

OpenStudy (donka74):

3 feet

OpenStudy (donka74):

if a yard is 98$ u only can get 3 feet

OpenStudy (ikram002p):

 1088.9/98=1.11 cubic card that is coverd \[4\times\frac{ 1 }{ 6 }\times x = 1.11\] \[x= \frac{ 1.11\times 6 }{ 4 } =1.665 yards = 1.665 \times 3 feets =4.995 feet\] so the answer would be b which is 5 :)